# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v1

@@ 3.2 Data generation of microscopic distributions
We generate synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy. Each simulation has a time span of $275\,\text{s}$. To reduce computational storage costs while retaining essential temporal features, we subsample the resulting pedestrian trajectories at intervals of $\delta t=0.25\,\text{s}$. As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios. In particular, we consider $C=10$ different initializations with zero velocities and position distributions including: (i) uniform distribution to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) piecewise linear distributions with higher densities near the wall boundaries to simulate characteristic wall-avoidance behavior, (iv) sinusoidal distributions to generate periodic macroscopic behavior, and (v) double Gaussian distributions to capture the formation of two distinct crowds. Details on the form and parameters of the above initializations are provided in Appendix B, distinguishing between those used for training and testing the proposed framework in Tables B.1 and B.2, respectively. The resulting datasets are comprised of 10 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1100}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1100$ timesteps are sampled from each trajectory. Thus, the total datasets overall cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(10\cdot 1100)}$ used for training and testing, respectively.
@@ 3.3 Extraction of continuous density fields
For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1. To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behaviors, while minimizing artificial grid effects. For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq. (6) tuned with bandwidths $\mathbf{H}=diag(h_{x},h_{y})=(3,2)$ for suppressing noise and preserving features of the microscopic data, while ensuring that local structures remain clear without introducing spurious artifacts. To account for the physical obstacle in $\Omega$, a binary mask is applied to explicitly zero out density estimation on the cell centroids of the obstacle geometry. To account for the periodic boundary conditions, we further perform domain augmentation for the KDE support. All the details regarding KDE implementation are provided in Appendix C.
Using the above procedure, we derive the density fields $\rho^{(c)}\big{(}x_{i},y_{j},t_{k}\big{)}$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing data sets $\mathcal{X}_{ts}$. For conforming to the assumptions of Proposition 1, we finally normalize each density field following Eq. (19), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq. (8), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=10\cdot 1100$.
@@ 3.4 POD for the restriction and lifting operators
3.5 Autoregressive ROMs and their training
@@
Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct autoregressive ROMs for learning the dynamics in the latent space. For this purpose, we learn linear MVARs and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$. To conform to this sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq. (7) on the training data $X_{tr}$.
@@
As a preliminary step, we first tune the lag window ww–within a bound of $w_{\text{max}}=100$–using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [88, 89, 90], the details of which are described in Appendix D. Both criteria are likelihood-based, with AIC tending to select larger windows than BIC. We note that BIC and AIC are well-suited for MVARs models since they assume stationarity of the data and white noise residuals. For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
@@
For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey-Fuller (ADF) test [91] with a threshold of significance level $p<0.01$. Then, we solve the least-squares optimization problem in Eq. (24) using ridge regression with a regularization parameter $\lambda=10^{-6}$ to compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq. (22).
@@
For the LSTM model, we consider a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq. (28). To train the LSTM model, we solve the optimization problem in Eq. (29) using the Adam optimizer [86] with automatic differentiation for the gradients of the loss function, enabled through MATLAB’s Deep Learning Toolbox [92]. The model’s parameters are initialized with a Glorot uniform initialization [93], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints. The maximum number of epochs was set to $100$ as no further improvement was achieved using more epochs. To further avoid overfitting of the LSTM, we split the training data into training and validation sets using a 80-20% ratio, and shuffle each data set without breaking the internal temporal structure of the sequences. We compute the loss function value on the validation set for every epoch and stop the training after 5 consecutive validation checks with no improvement.
To assess the convergence of both MVARs and LSTMs training, we report the values of the loss functions in Eqs. (25) and (30), respectively, at the end of training. Since the MVARs loss function is based on $L_{2}$ error, we further report the MSE:
@@
where $\hat{\boldsymbol{y}}_{c}(t_{k})$ is the prediction in Eq. (26), to enable direct comparison with the MSE-based loss function of LSTMs. Additionally, we report the computational times required for training both ROMs; we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the ridge regression employed for solving the related optimization problem.
@@ 3.6 Accuracy assessment of the proposed framework
Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $c=10$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details. In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$. Then, following the proposed restrict-evolve with ROM-lift framework in Eq. (5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities; i.e., computing:
@@
where $\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators in Eqs. (7) and (21) and $\Phi_{\text{ROM}}$ is the autoregressive ROM with parameters $\mathbf{p}$ that can be either the trained MVARs or LSTMs model. Notice that while Eq. (35) could consider the restricted values of the normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, it instead considers the predicted normalized densities; thus, evaluating the proposed framework in a closed-loop (autoregressive) manner. This allows us to feed subsequent predictions recursively into Eq. (35) to advance the forecast over time.
To quantify the prediction accuracy, we compute the relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ per case $c$ and time step $k$ as:
@@
We report these errors overall cases and time steps, also computing the 10-90 percentiles, and track down the evolution of these errors along time steps.
@@ 4 Numerical results
At the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3. Next, we employed POD on $\mathcal{X}_{tr}$ to discover the latent space dimension; see Section 3.4. Following the criterion in Eq. (32), the minimum latent dimension required for retaining more than 99% of the energy in the dataset is $d=13$ modes, as shown in Fig. 3a).
@@
Figure 3: Accuracy of the restriction and lifting operators in Eqs. (7) and (21). Panel a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set. Panel b) depicts the average (over the $c=10$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq. (33) with $d=13$ POD modes, in time over the training and testing datasets.
@@
Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs. (7) and (21). To assess their accuracy, we computed the relative $L_{2}$ reconstruction error $e^{2,rec}_{k}$ in Eq. (33) across the $k=1,\ldots,K$ time steps of each trajectory. Figure 3b) displays the average relative error overall $c=10$ cases with different initial conditions included in either the training or testing datasets. While in the beginning, the reconstruction error is more than 10%, after the first $100$ time steps, it reduces to 5% for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$. We note that this level of reconstruction error serves as a baseline for the proposed framework, since the timeseries forecasting with ROMs in the latent space introduces additional approximation errors.
@@
Having constructed the restriction and lifting operators (second and fourth step of the framework), we next train MVARs and LSTMs autoregressive ROMs to learn the latent dynamics; see Section 2.2.3. First, we obtained the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ from the density profiles of the training set, using the restriction operator. We then determined the lag window, using the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas the AIC favored (a wider, as expected) $w=9$. Both windows were used for training both MVARs and LSTMs models, denoted from now on as MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models. For performing sequence-to-one learning, we further divided the restricted training data set into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
@@
For the MVARs training, we first verified that the latent dynamics datasets $Y_{f},Y_{t}$ constitute stationary time series. Using the ADF test, we confirmed that all $d$ variables are stationary processes with a significance level of $p=0.001$. We next trained all MVAR and LSTM models using algorithms, hyperparameters and stopping criteria discussed in detail in Section 3.5. The training results are shown in Table 1 for the four ROMs, including the loss function values attained at the end of training and the computational time (in seconds) required.
@@
Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models. Loss function values at the end of training for MVARs $\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{c})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)$ in Eq. (25) and for LSTMs $\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{c})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)$ in Eq. (30), along with computational times (in seconds) required for training. For MVARs, the MSE in Eq. (34) is reported for comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
@@
Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s) $\downarrow$
MVAR(4) $1.66\times 10^{-3}$ $1.96\times 10^{-11}$ - 0.0055
MVAR(9) $1.01\times 10^{-3}$ $7.38\times 10^{-12}$ - 0.042
LSTM(4) - - $8.64\times 10^{-7}$ 180
LSTM(9) - - $3.51\times 10^{-7}$ 443
@@
Since the MVARs loss function is $L_{2}$ based, we further report the MSE, computed in Eq. (34), for MVARs to allow comparison with the MSE-based loss of the LSTMs. Table 1 indicates that all the optimization problems converge, as evident from the low loss function values attained. MVARs are able to reach much smaller loss function values than LSTMs, providing a hint for their better accuracy. As expected, LSTMs require significantly higher computational time for training in comparison to MVARs. Table 1 further shows that the training of both types of ROM converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
@@
Having constructed and trained all components of our framework, we now evaluate its end-to-end performance, using the forecasting closed-loop time-stepper in Eq. (35) for long-horizon forecasting. In particular, we quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ for the $c$, unseen in training, cases–of different initial conditions–included in the testing set $X_{ts}$ across $k$ time steps. Table 2 provides a summary of the relative errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ for all four ROMs, including the median and the 10-90 percentiles, overall cases and time steps of the testing set.
@@
Table 2: End-to-end numerical accuracy of the closed-loop framework across the testing set, using the trained MVARs and LSTMs models for the latent dynamics. The relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors of the high-dimensional density space, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) are reported, for the four MVAR(4), MVAR(9), LSTM(4), LSTM(9) latent dynamics models; median and 10-90 percentiles (in parenthesis) are shown overall $C=10$ cases and $K=1091$ time steps.

### v2

@@ 3.2 Data generation of microscopic distributions
We generate synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy. Each simulation has a time span of $250\,\text{s}$. To reduce computational storage costs while retaining essential temporal features, we subsample the resulting pedestrian trajectories at intervals of $\delta t=0.25\,\text{s}$. As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios. In particular, we consider $C=20$ different initializations with zero velocities and positions distributed via: (i) uniform distributions to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) “double bell-shaped” Gaussian distributions to capture the formation of two distinct crowd groups and (iv) sinusoidal distributions to generate periodic macroscopic behavior. Details on the form and parameters of the above initializations are provided in Appendix B, distinguishing between those used for training and testing the proposed framework in Tables B.1 and B.2, respectively. The resulting datasets are comprised of 20 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1000}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1000$ time steps correspond to the samples recorded from each trajectory. Thus, the total datasets over all cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(20\cdot 1000)}$ used for training and testing, respectively.
@@ 3.3 Extraction of continuous density fields
For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1. To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behavior, while minimizing artificial grid effects. For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq. (6) tuned with bandwidths $\mathbf{H}=diag(h_{x},h_{y})=(3,2)$ for suppressing noise and preserving features of the microscopic data, while ensuring that local structures remain clear without introducing spurious artifacts. To account for the physical obstacle in $\Omega$, a binary mask is applied to explicitly zero out density estimation on the cell centroids of the obstacle geometry. To account for the periodic boundary conditions, we further perform domain augmentation for the KDE support. All the details regarding KDE implementation are provided in Appendix C.
Using the above procedure, we derive the density fields $\rho^{(c)}\big(x_{i},y_{j},t_{k}\big)$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing $\mathcal{X}_{ts}$ data sets. For conforming to the assumptions of Proposition 1, we finally normalize each density field following Eq. (19), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq. (9), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=20\cdot 1000$.
@@ 3.4 POD for the restriction and lifting operators
3.5 ROMs for latent space dynamics and their training
@@
Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct ROMs for learning the dynamics in the latent space. For this purpose, we learn linear MVAR and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$. To conform to this sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq. (20) on the training data $X_{tr}$.
@@
As a preliminary step, we first tune the lag window $w$, within a bound of $w_{\text{max}}=100$, using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [93, 94, 95]; details are provided in Appendix D. Both criteria are likelihood-based, with AIC tending to select larger windows than BIC. We note that BIC and AIC are well-suited for MVARs models since they assume stationarity of the data and white noise residuals. For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
@@
For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey–Fuller (ADF) test [96] with a threshold of significance level $p<0.01$. Then, we compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq. (22) using ordinary least-squares (OLS) to solve the optimization problem in Eq. (24).
@@
For the LSTM model, we consider a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq. (28). To train the LSTM model, we solve the optimization problem in Eq. (29) using the Adam optimizer [91] with automatic differentiation for the gradients of the loss function, enabled through MATLAB’s Deep Learning Toolbox [97]. The model’s parameters are initialized with a Glorot uniform initialization [98], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints. The maximum number of epochs was set to $200$ as no further improvement was achieved using more epochs. To further avoid overfitting of the LSTM, we split the training data into training and validation sets using a 80-20% ratio, and shuffle each data set without breaking the internal temporal structure of the sequences. We compute the loss function value on the validation set for every epoch and stop the training after 5 consecutive validation checks with no improvement.
To assess the convergence of both MVARs and LSTMs training, we report the values of the loss functions in Eqs. (25) and (30), respectively, at the end of training. To further quantify the uncertainty associated with the stochastic training of the LSTM models, we repeated their training procedure 50 times, each with different randomly initialized learnable parameters. We report the mean error of the final losses and the 10–90% error percentiles across the 50 runs. Since the MVARs loss function is based on $L_{2}$ error, we further report the MSE at the latent space:
@@
where $\hat{\boldsymbol{y}}_{c}(t_{k})$ is the prediction in Eq. (26), to enable direct comparison with the MSE-based loss function of LSTMs. Additionally, we report the computational times required for training both ROMs (for LSTMs, we report the mean error and 10–90% error percentiles statistics across the 50 runs); we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the OLS method employed for solving the related optimization problem.
@@ 3.6 Accuracy assessment of the proposed framework
Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $C=20$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details. In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$, using Eq. (19). Then, following the proposed restrict-evolve with ROM-lift framework in Eq. (5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities in both an open-loop and closed-loop setting. For the open-loop setting evaluation, we compute the prediction:
$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(x^{(c)}(t_{k-1})),\ldots,\mathcal{R}(x^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (35)
while for the closed-loop setting evaluation, we compute the prediction:
@@
where $\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators in Eqs. (20) and (21) and $\Phi_{\text{ROM}}$ is the latent dynamics ROM with parameters $\mathbf{p}$ that can be either the trained MVAR or LSTM model. In particular, for the LSTM model, we consider the parameter set resulting in the best training performance among the 50 runs. Notice that the open-loop prediction in Eq. (35) considers the restricted values of the observed normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, while the closed-loop prediction in Eq. (36) instead considers the predicted–from previous steps–normalized densities $\mathcal{R}(\hat{x}^{(c)}(t_{k-j}))$ in a recursive fashion. Since the open-loop setting requires the “ground truth” densities, we use it here to evaluate the one-step predictions of our framework. For the desired, in practice, longer predictions, we utilize the closed-loop setting, which enables us to recursively feed subsequent predictions into Eq. (36), resulting in an autoregressive time-stepping framework also considering the effect of one-step forecast error accumulation.
To quantify the end-to-end prediction accuracy, we compute the relative, reconstructed in the ambient-space, $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ (either for the open-loop or the closed-loop setting) per case $c$ in the testing set and time step $k$ as:
@@
For a summary of statistics, we report the mean and the 10–90% percentiles of the above errors over all cases and time steps. To track down the prediction accuracy in time, we further depict the mean and 10–90% percentiles of the error evolution in time over all cases.
@@ 4 Numerical results
In the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3. Next, we employed POD on $\mathcal{X}_{tr}$ to discover the latent space dimension; see Section 3.4. Following the criterion in Eq. (32), the minimum latent dimension required for retaining more than 99% of the energy in the training dataset is $d=13$ modes, as shown in Fig. 3(a).
@@
Figure 3: Accuracy of the restriction and lifting operators in Eqs. (20) and (21). Panel (a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set. Panel (b) depicts the average (over the $C=20$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq. (33) with $d=13$ POD modes, in time for the training and testing datasets.
@@
Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs. (20) and (21). To assess their accuracy, we computed the relative $L_{2}$ reconstruction error $e^{2,rec}_{k}$ in Eq. (33) across the $k=1,\ldots,K$ time steps of each trajectory. Fig. 3(b) displays the average relative error over all the $C=20$ cases with different initial conditions included in either the training or testing datasets. While initially the reconstruction error is $\sim$15%, it decreases to 10% at the first 30s30\ $\text${s} and further reduces to $\sim$7% after 50s50\ $\text${s} for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$. We note that this level of reconstruction error serves as an infimum approximation error baseline, i.e., the best possible error for the proposed framework at the ambient–density profiles–space. ROM forecasts in the long term are therefore expected to surpass this baseline error due to the additional error contributions from model-form error and training/generalization errors.
@@
We next train MVAR and LSTM models to learn ROMs for the latent dynamics (see Section 2.2.3), where the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ are obtained from the density profiles of the training set via the restriction operator. To determine the lag window, we employed the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas AIC favored a wider (as expected) $w=9$. Both windows were used for training both MVAR and LSTM models, denoted from now on as MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models. The training data sets were divided into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
@@
For the MVARs training, we first verified that the latent dynamics datasets $Y_{f},Y_{t}$ constitute stationary time series; the ADF test confirmed stationarity with a significance level of $p<0.001$ for all $d$ variables. We next trained all MVAR and LSTM models using algorithms, hyperparameters and stopping criteria discussed in detail in Section 3.5. The training results are shown in Table 1 for the four ROMs, including the loss function values attained at the end of training and the computational time (in seconds) required. For the LSTMs, as each model was trained over 50 runs with different parameter initializations, we report in Table 1 the mean and the 10–90% error percentiles across the 50 runs.
@@
Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models at the latent space. Loss functions $\mathcal{L}_{\text{MVAR}}\left(\cdot\right)$ in Eq. (25) for MVARs and $\mathcal{L}_{\text{LSTM}}\left(\cdot\right)$ in Eq. (30) for LSTMs at the end of training, along with computational times (in seconds) required for training. For the LSTMs, the mean and 10–90% error percentiles (in parentheses) are reported across the 50 training processes with different parameter initializations. Additionally, the MSE loss function of MVARs in Eq. (34) is reported for comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
@@
Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s)
MVAR(4) $1.91\times 10^{-4}$ $1.41\times 10^{-13}$ - $0.0021$
MVAR(9) $1.34\times 10^{-4}$ $7.05\times 10^{-14}$ - $0.0045$
LSTM(4) - - $2.35\ (2.22,2.50)\times 10^{-5}$ 252(236,302)252\ (236,302)
LSTM(9) - - $1.81\ (1.68,1.93)\times 10^{-5}$ 310(302,326)310\ (302,326)
@@
Since the loss function of MVARs is $L_{2}$-based, we additionally compute the MSE in Eq. (34) to enable direct comparison with the MSE-based LSTM loss. MVARs are able to reach much smaller loss function values than LSTMs, providing a hint for their better accuracy. As expected, LSTMs require significantly higher computational time for training in comparison to MVARs. Table 1 further shows that the training of both types of ROMs converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
One-Step/Open-loop Predictions
@@
Having constructed and trained all components of our framework, we now evaluate its end-to-end performance in one-step predictions in the ambient–density profile–space using the open-loop time-stepper in Eq. (35); note that for the LSTMs, we consider the model with the best training performance among the 50 model runs. We quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$, and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (37), computed between the “ground-truth” density profiles $x^{(c)}(t_{k})$ and the predicted/reconstructed ones $\hat{x}^{(c)}(t_{k})$ for the $C$, unseen in training, cases (of different initial conditions) included in the testing set $X_{ts}$ across time steps $k$. Table 2 provides a summary of the relative reconstructed errors in the ambient space, for one-step predictions, for all four ROMs, including the mean error and the 10–90% error percentiles, aggregated over all cases and time steps in the testing set.
@@
Table 2: Open-loop (one-step) prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models. The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models. Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).